OpenStudy (anonymous):

If h(x)=(fog)(x) and h(x)=3(x+2)^2, find f(x) and g(x).

OpenStudy (raden):

one of f(x) and g(x) must be knowed, right ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

so, why didnt you write it ?

OpenStudy (raden):

ohh, by using trial error one by one ...

OpenStudy (anonymous):

yea -.- i couldnt seem to get it

OpenStudy (raden):

if f(x) = x+2 and g(x) = 3x^2 then (fog)(x) = f(g(x)) = f(3x^2) = 3x^2 + 2, so it is wrong because we hope get (fog) = 3(x+2)^2

OpenStudy (raden):

can you check for the 3rd one :)

OpenStudy (raden):

or maybe the last one be true ?

OpenStudy (anonymous):

mkay hold on

OpenStudy (raden):

sorry, i have to now. i sure @wio can help you :) byyeeee

OpenStudy (raden):

* go